OpenStudy (kamille):

if lg5=a, lg3=b when \[\lg _{30}8\] is? By the way, lg stands for log10

OpenStudy (kamille):

how i think: \[\log _{30}8=\frac{ \lg8 }{ \lg30 }=\frac{ \lg24-\lg3}{ \lg10+\lg3 }=\frac{ \lg12+\lg2-b }{ \lg2+a+b }=\] and I am not sure what to do nextt

OpenStudy (raden):

hmm.. looks u need the value of lg 2 (in term of a)

OpenStudy (raden):

can u ?

OpenStudy (kamille):

oh wrong

OpenStudy (kamille):

can you show me?

OpenStudy (raden):

not wrong, but just need extra important point there, is the value of log 2

OpenStudy (raden):

look this : log 2 = log (10/5) = log 10 - log 5 = 1 - a now, apply it into ur equation above

OpenStudy (raden):

btw, that will be easier if u setting log 8 = log (2^3) = 3 log 2

OpenStudy (raden):

and log 30 = log (2 * 3 * 5) = log 2 + log 3 + log 5

OpenStudy (kamille):

well, log5 isnt the same as lg5

OpenStudy (kamille):

so I still didnt get it, sorry

OpenStudy (raden):

log 8/log 30 = log (2^3)/log(2*3*5) = 3 log2/(log2+log3+log5) = 3 (1-a)/(1-a+b+a) = .... simplify

OpenStudy (raden):

got it ?

OpenStudy (kamille):

thanks

OpenStudy (kamille):

well, i will analyze;)